OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.04x + 3.14 = 7.98 A. x ≈ 4.7 B. x ≈ 10.7 C. x ≈ 4.6 D. x ≈ 4.5

OpenStudy (anonymous):

1.04x + 3.14 = 7.98 Isolate the x term, so we move the 3.14 to the other side of the equal sign 1.04x = 7.98 - 3.14 simpify 1.04x = 4.84 isolate x by dividing both sides by 1.04 x = 4.84/1.04 simplify the fraction with a calculator x = 4.65384 then round to the nearest tenth x = 4.7

OpenStudy (anonymous):

hey, lets not give answers, it doen'st teach usually..

OpenStudy (anonymous):

but keep on helping people, just not answer giving please ^_^

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks :) @DemolisionWolf

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

thank you both :DD sooo much ^^